HIGH RESOLUTION SPECTROSCOPY OF ZrN: THE (0,0) BAND OF THE A^{2}Π − X^{2}Σ^{+} TRANSITION \underline{A. S-C. CHEUNG}, HAIYANG LI, C. M-T. CHAN and HONGBING CHEN, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.